a battery of 10V is connected with 5 ohm resistance then find current in this circuit.
1
Expert's answer
2020-09-21T08:33:08-0400

According to the Ohm's law,

"\displaystyle V= IR \Rightarrow \; I = \frac{V}{R} = \frac{10 \, V}{5 \, \Omega} = 2\, A."

Answer: 2 A.